\begin{thebibliography}{10}


\bibitem{Heise} 
Heisenberg, W., Mehrkoerperproblem und Resonanz in der Quanten-mechanik. Zeitschrift fur Physik 38, 411-426 (1926). 
\bibitem{fermi}
Fermi, E., (1926). Sulla quantizzazione del gas perfetto monoatomico. Rendiconti della Accademia dei Lincei 3:145-149.
\bibitem{Dirac}
Dirac P. A. M., (1926). On the Theory of Quantum Mechanics. Proceedings of the Royal Society A 112:661-677.
\bibitem{Wentzel}
Wentzel, G., (1927). Ueber strahlungslose Quantensprünge. Zeitschrift fur Physik, 43:524-530.  
%doi:10.1007/BF01397631.
\bibitem{majorana} 
Majorana, E., (1931). Teoria dei tripletti incompleti. Il Nuovo Cimento (1924-1942) 8:107.  
%doi: 10.1007/BF02959482
\bibitem{Foote} 
Foote, P. D., Takamine, T., Chenault, R. L., (1925). The Excitation of Forbidden Spectral Lines.  Phys. Rev., 26:165. 
%doi:10.1103/PhysRev.26.165
\bibitem{Shenstone} 
Shenstone, A. G. (1931). Ultra-Ionization Potentials in Mercury Vapor. Phys. Rev., 38:873
%doi: 10.1103/PhysRev.38.873
\bibitem{Beutler} 
Beutler, H., (1935). Über Absorptionsserien von Argon, Krypton und Xenon zu Termen zwischen den beiden Ionisierungsgrenzen $^2P_3$ $_{2/0}$ und $^2P_1$ $_{2/0}$ Zeitschrift für Physik, 93:177-196.
\bibitem{Fano} 
Fano U. (1935). On the absorption spectrum of noble gases at the arc spectrum limit. l Nuovo Cimento (1924-1942) 12:154-161
\bibitem{Blatt1}
Blatt, J. M., Weisskopf V.F. (1952) Theoretical nuclear physics, J. Wiley and sons, New York
\bibitem{Blatt2} 
Blatt, J. M., Weisskopf V. F., Critchfield C. L., (1953). Theoretical Nuclear Physics Am. J. Phys. 21:235
doi: 10.1119/1.1933407 
\bibitem{Fes1}  
Feshbach H.,  Porter C. E.,  Weisskopf V. F. (1954). Model for Nuclear Reactions with Neutrons Phys. Rev. 96:448
\bibitem{Fes2} 
Feshbach H. (1958). Unified theory of nuclear reactions, Annals of Physics 5, 357-390
\bibitem{Vittorini}
 Vittorini-Orgeas, A., Bianconi, A., (2009) From Majorana Theory of Atomic Autoionization to Feshbach Resonances in High Temperature Superconductors. Journal of Superconductivity and Novel Magnetism, 22:215-221
doi:10.1007/s10948-008-0433
\bibitem{Fano2} 
Fano, U. (1961). Effects of Configuration Interaction on Intensities and Phase Shifts. Phys. Rev., 124:1866-1878
doi: 10.1103/PhysRev.124.1866
\bibitem{Madden1} 
Madden, R. P. and Codling, K. (1963). 
New autoionizing atomic energy levels in He, Ne, and Ar. Physical Review Letters, 10(12):516-518.
\bibitem{Madden2} 
Madden, R. P. and Codling, K. (1965). 
Two-Electron excitation states in helium. The Astrophysical Journal, 141:364.
\bibitem{Cooper} Cooper, J. W., Fano, U., and Prats, F. (1963). 
Classification of Two-Electron excitation levels of helium. Physical Review Letters, 10(12):518-521.
\bibitem{Anderson}
Anderson P. W. (1961) Localized Magnetic States in Metals Phys. Rev. 124:41
 %doi:10.1103/PhysRev.124.41
\bibitem{Kondo}
Kondo, J. (1964). Resistance Minimum in Dilute Magnetic Alloys". Progress of Theoretical Physics 32: 37
 %doi:10.1143/PTP.32.37.
\bibitem{dhez} Dhez P. and Ederer D. L. (1973) Photoionization resonance profile parameters of two-electron excitation in He. Journal of Physics B: Atomic and Molecular Physics, 6(4):L59-L64.  43(25):1892-1896.
  \bibitem{bernieri}
  Bernieri, E. and Burattini, E. (1987) Multielectron transitions above the krypton K edge. Physical Review A, 35(8):3322-3326.
\bibitem{soldatov}
   Soldatov, A. V., Ivanchenko, T. S., Della Longa, S., and Bianconi, A. (1993). 
Two-electron excitations and one-electron multiple-scattering resonances in the x-ray absorption of solid neon. Physical Review B, 47(24):16155-16161.
 \bibitem{bianconi1991a}
Bianconi, A., Garcia, J., Benfatto, M., Marcelli, A., Natoli, C. R., and Ruiz-Lopez, M. F. (1991) 
Multielectron excitations in the K -edge X-ray-absorption near-edge spectra of V, Cr, and
 Mn 3$d^0$ compounds with tetrahedral coordination. Physical Review B, 43(9):6885-6892.
 \bibitem{dicicco}
   Di Cicco, A., Stizza, S., Filipponi, A., Boscherini, F., and Mobilio, S. (1992).
 X-ray absorption investigation of Si$X_4$ ($X$=Cl, F,  C$H_3$). Journal of Physics B: Atomic, Molecular and Optical Physics, 25(10):2309-2318.
 \bibitem{dangelo}
 DAngelo, P., Petit, P.-E., and Pavel, N. V. (2004) Double-Electron excitation channels at the Ca K-Edge of hydrated calcium ion. J. Phys. Chem. B, 108(31):11857-11865.
\bibitem{chaboy1990}
  Chaboy, J., Garcia, J., Marcelli, A. and Ruiz Lopez, M.F. (1990). 
An experimental and theoretical study of multielectron excitations at the L3 absorption edge in some rare earth alloys and their hydrides'
Chem. Phys. Lett. 174:389
\bibitem{chaboy1994}
Chaboy, J., Marcelli, A. and Tyson, T.A. (1994) 
Influence of double-electron transitions on the EXAFS L edges of rare-earth systems.
Phys. Rev. B 49:11652-11661
\bibitem{roulet}
Roulet, B., Gavoret, J., and Nozières, P. (1969). 
Singularities in the X-Ray absorption and emission of metals. i. First-Order parquet calculation. Physical Review, 178(3):1072-1083.
\bibitem{nozieres}
Nozieres, P. and De Dominicis, C. T. (1969). 
Singularities in the X-Ray absorption and emission of metals. III. One-Body theory exact solution. Physical Review, 178(3):1097-1107.
\bibitem{shotte}
Schotte, K. D. and Schotte, U. (1969). 
Tomonaga's model and the threshold singularity of X-ray spectra of metals. Physical Review, 182(2):479-482.
\bibitem{mahan} 
Mahan, G. D. (1971). 
Excitonic effects in x-ray transitions in metals. edited by Bennett, L. H., editor, Electronic Density of States, 3rd Materials Research Symposium,Gaithersburg, Maryland Nov 3-8, 1969, pages 253-259. National Bureau of Standards, Washington, D.C.: U.S. Gov. Print. Off..
\bibitem{heidin} 
Hedin, L., Lundqvist, B., and Lundqvist, S. (1971). 
Beyond the one-electron approximation : Density of states for interacting electrons. In Bennett, L. H., editor, Electronic Density of States  3rd 
Materials Research Symposium, Gaithersburg, Maryland Nov 3-8, 1969, pages 233-248. National Bureau of Standards, Washington, D.C.: U.S. Gov. Print. Off., 
\bibitem{combescot}
Combescot, M. and Nozieres, P. (1971). 
Infrared catastrophy and excitons in the x-ray spectra of metals. Journal de Physique, 32(11-12):913-929.
\bibitem{doniach}
Doniach, S. and Sunjic, M. (1970). Many-electron singularity in x-ray photoemission and x-ray line spectra from metals. Journal of Physics C: Solid State Physics, 3(2):285-291.
\bibitem{antonangeli}
Antonangeli, F., Balzarotti, A., Bianconi, A., Perfetti, P., Ascarelli, P., and Nisticò, N. (1977). 
Many-body contribution to XPS line asymmetry in palladium-hydrogen systems. Il Nuovo Cimento B (1971-1996), 39(2):720-726.
\bibitem{benfatto1983}
Benfatto, M., Bianconi, A., Davoli, I., Incoccia, L., Mobilio, S., and Stizza, S. (1983). 
Role of multielectron excitations in the $L_3$ XANES of Pd. Solid State Communications, 46(5):367-370.
\bibitem{bachrach1982}
Bachrach, R. Z. and Bianconi, A. (1982). 
Carbon K shell near threshold photoemission cross-section: Energy dependence of elastic peak and plasmon sideband in graphite. Solid State Communications, 
42(7):529-532
\bibitem{balz1974}
 Balzarotti, A., Bianconi, A., and Burattini, E. (1974). 
Role of the density of conduction states on the $L_2,$$_3$ absorption spectrum of aluminum. Physical Review B, 9(12):5003-5007.
 \bibitem{balza1974}
Balzarotti, A., Bianconi, A., Burattini, E., Grandolfo, M., Habel, R., and Piacentini, M. (1974). 
Core transitions from the al 2p level in amorphous and crystalline $Al_2O_3$. Physica Status Solidi (b), 63(1):77-87.
\bibitem{BianconiPetersen1978}
Bianconi, A., Petersen, H., Brown, F., and Bachrach, R. Z. (1978). 
K-shell photoabsorption spectra of $N_2$ and $N_2O$ using synchrotron radiation. Physical Review A, 17(6):1907-1911.
\bibitem{Langhoff974}
Langhoff, S. R. and Davidson, E. R. (1974). 
Configuration interaction calculations on the nitrogen molecule. International Journal of Quantum Chemistry, 8(1):61-72
\bibitem{Broad1976}
Broad, J. T. and Reinhardt, W. P. (1976). 
One- and two-electron photoejection from $H^-$: A multichannel j -matrix calculation. Physical Review A, 14(6):2159-2173.
\bibitem{Dehmer1975}
Dehmer, J. L. and Dill, D. (1975). 
Shape resonances in K-shell photoionization of diatomic molecules. Physical Review Letters, 35(4):213-215.
\bibitem{Dehmer1976}
Dehmer, J. L. and Dill, D. (1976). 
Molecular effects on inner?shell photoabsorption. K shell spectrum of $N_2$. The Journal of Chemical Physics, 65(12):5327-5334.
\bibitem{Dehmer1979}
Dehmer, J. L., Dill, D., and Wallace, S. (1979). 
Shape-Resonance-enhanced Nuclear-Motion effects in molecular photoionization. Physical Review Letters, 43(14):1005-1008.
\bibitem{Loomba1981}
Loomba, D., Wallace, S., Dill, D., and Dehmer, J. L. (1981). 
Pictures of unbound molecular electrons, including shape resonant states: eigenchannel contour maps. The Journal of Chemical Physics, 75(9):4546-4552.
\bibitem{BianconiDoniach1978}
Bianconi, A., Doniach, S., and Lublin, D. (1978). 
X-ray Ca K edge of calcium adenosine triphosphate system and of simple Ca compounds. Chemical Physics Letters, 59(1):121-124.
\bibitem{Bianconi1979}
Bianconi, A. (1979).
 Core excitons and inner well resonances in surface soft x-ray absorption (SSXA) spectra. Surface Science, 89(1-3):41-50.
\bibitem{Bianconi1980}
Bianconi, A. (1980). Surface X-Ray absorption spectroscopy: Surface EXAFS and surface XANES. Applications of Surface Science, 63(3-4):392-418.
\bibitem{Bianconi1982}
Bianconi, A., Dell'Ariccia, M., Durham, P., and Pendry, J. (1982). 
Multiple-scattering resonances and structural effects in the x-ray-absorption near-edge spectra of Fe II and Fe III hexacyanide complexes. Physical Review B, 26(12):6502-6508.
\bibitem{DellaLonga1995}
Della Longa, S., Soldatov, A., Pompa, Bianconi, A., (1995) Atomic and electronic structure probed by x-ray absorption spectroscopy: Full multiple scattering analysis with the G4XANES package. Computational Materials Science 4 (3):199-210.
% http://dx.doi.org/10.1016/0927-0256(95)00027-n
\bibitem{garciaBianconi1986}
 Garcia, J., Bianconi, A., Benfatto, M., Natoli, C. R., (1986) Coordination geometry of transition metal ions in dilute solutions by XANES. Le Journal de Physique Colloques 47 (C8):C8-49-C8-54.
%URL http://dx.doi.org/10.1051/jphyscol:1986807
\bibitem{garcia1986}
 Garcia, J., Benfatto, M., Natoli, C.R.,Bianconi, A., Davoli I., and Marcelli A. (1986)
Three particle correlation function of metal ions in tetrahedral coordination determined by XANES
Solid State Commun. 58:595-599 
\bibitem{Rao1980}
  Rao, C. N. R., Sarma, D. D., Sarode, P. R., Sampathkumaran, E. V., Gupta, L. C., and Vijayaraghavan, R. (1980). 
Valence fluctuation in some Yb intermetallics by x-ray photoemission and x-ray absorption. Chemical Physics Letters, 76(3):413-415.
\bibitem{Launois1980}
 Launois, H., Rawiso, M., Holland-Moritz, E., Pott, R., and Wohlleben, D. (1980). 
X-Ray absorption study of Mixed-Valence TmSe. Physical Review Letters, 44(19):1271-1275.
\bibitem{BianconiModesti1981} 
  Bianconi, A., Modesti, S., Campagna, M., Fischer, K., and Stizza, S. (1981). 
X-ray absorption near edge structures of intermediate valence TmSe single crystals. Journal of Physics C: Solid State Physics, 14(31):4737-4743.
\bibitem{Sarode1981} 
Sarode, P. R., Sarma, D. D., Rao, C. N. R., Sampathkumaran, E. V., Gupta, L. C., and Vijayaraghavan, R. (1981). 
X-ray absorption spectroscopic study of the mixed valence system Ce$Pd_3$. Materials Research Bulletin, 16(2):175-178.
\bibitem{Steglich1979}
Steglich, F., Aarts, J., Bredl, C. D., Lieke, W., Meschede, D., Franz, W., Schäfer, H. (1979). Superconductivity in the presence of strong Pauli paramagnetism: $Ce Cu_2Si_2$ 
Phys. Rev. Lett. 43: 1892
\bibitem{Jarlborg1983}
Jarlborg, T., Braun, H. F., and Peter, M. (1983). 
Structural properties and band structure of heavy fermion systems: $CeCu_2Si_2$ and $LaCu_2Si_2$. 
Zeitschrift für Physik B Condensed Matter 52: 295-301.
\bibitem{Lieke1982} 
Lieke, W., Rauchschwalbe, U., Bredl, C. B., Steglich, F., Aarts, J., and de Boer, F. R. (1982). 
Superconductivity in $CeCu_2Si_2$. Journal of Applied Physics, 53(3):2111-2116.
\bibitem{BianconiCampagna1981} 
Bianconi, A., Campagna, M., Stizza, S., and Davoli, I. (1981). 
Intermediate valence and near-edge structure in the x-ray absorption spectrum of CePd$_3$, $\gamma$ -Ce, and $CeCu_2Si_2$. Physical Review B, 24(10):6139-6142.
\bibitem{Padalia1981}
Nagarajan, R., Sampathkumaran, E. V., Gupta, L. C., Vijayaraghavan, R., Bhaktdarshan, and Padalia, B. D. (1981). 
X-ray absorption spectroscopic study of a mixed valence system, $EuPd_2Si_2$. Physics Letters A, 81(7):397-398.
\bibitem{BianconiCampagna1982}  
Bianconi, A., Campagna, M., and Stizza, S. (1982). Localization of screening orbitals, local-environment effects, and intermediate valence in core-level spectroscopy: Ce vs Tm. Physical Review B, 25(4):2477-2482.
\bibitem{Lengeler1983}  
Lengeler, B., Materlik, G., and Müller, J. E. (1983). 
L-edge x-ray absorption spectra of $\alpha$ and $\gamma$ cerium. Physical Review B, 28(4):2276-2278.
\bibitem{Raaen1983}
Raaen, S. and Parks, R. D. (1983). 
Mixed valence in $CeNi_5$; effects of dilution and chemical pressure. Solid State Communications, 48(2):199-202.
\bibitem{Parks1983} 
 Parks, R. D., Raaen, S., denBoer, M. L., Murgai, V., and Mihalisin, T. (1983). 
Anomalous saturation of mixed valence in cerium-based systems as studied by x-ray absorption. Physical Review B, 28(6):3556-3558.
\bibitem{BianconiMarcelli1984} 
Bianconi, A., Marcelli, A., Davoli, I., Stizza, S., and Campagna, M. (1984). 
Experimental evidence for the  shake-down peak in L III (and L II)-XANES of light rare earth intermetallics. Solid State Communications, 49(5):409-415.
\bibitem{Marcelli1984} 
Marcelli, A., Bianconi, A., Davoli, I., and Stizza, S. (1984) XANES of palladium rare earth intermetallics RPd3: determination of hybridization and mixing of 4f-orbitals. Proceedings International Conference EXAFS and Near Edge Structure III Stanford, 1984 eds. Hodgson, K.O., Hedman, B., and Penner-Hahn, J.E., Springer Science  Business Media, 11 Nov 2013, Springer Proceedings in Physics 2 (Springer) p.52
\bibitem{Marcelli1985} 
Marcelli, A., Bianconi, A., Davoli, I., and Stizza, S. (1985). 
Localization mixing and / or hybridization in intermetallic compounds RPd3 (R= La, Ce, Pr, Nd, Sm) by XANES. Journal of Magnetism and Magnetic Materials, 47-48:206-208.
\bibitem{BianconiMarcelli1985}
Bianconi, A. Marcelli, A., Tomellini, M., and Davoli, I. (1985)
Determination of mixing of 4f-ligand orbitals in Ce(SO4)2 by XANES. Is Ce(SO4)2 a mixed valent insulating system?.
in the Proceedings 4th International Conference on Valence Fluctuations (Köln, RFT 1984). J. Mag. Mag. Mat. 47-48: 209-211 
\bibitem{KotaniJo1987}  
Kotani, A., Jo, T., Okada, K., Nakano, T., Okada, M., Bianconi, A., Marcelli, A., and Parlebas, J. (1987). 
High energy spectroscopy in Ce and La compounds. Journal of Magnetism and Magnetic Materials, 70(1-3):28-32.
\bibitem{KotaniOkada1987} 
Kotani, A., Okada, M., Jo, T., Bianconi, A., Marcelli, A., and Parlebas (1987). 
Many body effect in inner shell photoemission and photoabsorption spectra of La compounds. Journal of the Physical Society of Japan, 56(2):798-809.
\bibitem{BianconiNatoli1978} 
 Bianconi, A. and Natoli, C. R. (1978).  Effect of the metal-insulator transition on vanadium K-photoabsortion spectrum in $V_2O_3$. Solid State Communications, 27(11):1177-1179.
\bibitem{Bianconi1982}
Bianconi, A. (1982). Multiplet splitting of final-state configurations in x-ray-absorption spectrum of metal $VO_2$: Effect of core-hole-screening, electron correlation, and metal-insulator transition. Physical Review B, 26(6):2741-2747.
\bibitem{Zaanen1985}
Zaanen, J., Sawatzky, G. A., and Allen, J. W. (1985). 
Band gaps and electronic structure of transition-metal compounds. Physical Review Letters, 55(4):418-421
\bibitem{Davoli1986}
Davoli, I., Marcelli, A., Bianconi, A., Tomellini, M., and Fanfoni, M. (1986). 
Multielectron configurations in the x-ray-absorption near-edge structure of NiO at the oxygen K threshold. Physical Review B, 33(4):2979-2982.
\bibitem{van der Laan1986}
van der Laan, G., Zaanen, J., Sawatzky, G. A., Karnatak, R., and Esteva, J. M. (1986). 
Comparison of x-ray absorption with x-ray photoemission of nickel dihalides and NiO. Physical Review B, 33(6):4253-4263.
\bibitem{Zaanen1986}
Zaanen, J., Westra, C., and Sawatzky, G. A. (1986). 
Determination of the electronic structure of transition-metal compounds: 2p x-ray photoemission spectroscopy of the nickel dihalides. Physical Review B, 33(12):8060-8073.
\bibitem{Petiau1986}
Petiau, J., Calas, G., Petitmaire, D., Bianconi, A., Benfatto, M., and Marcelli, A. (1986). 
Delocalized versus localized unoccupied 5f states and the uranium site structure in uranium oxides and glasses probed by x-ray-absorption near-edge structure. Physical review. B, Condensed matter, 34(10):7350-7361.
\bibitem{BianconiMarcelliDexpert1987}
Bianconi, A., Marcelli, A., Dexpert, H., Karnatak, R., Kotani, A., Jo, T., and Petiau, J. (1987). 
Specific intermediate-valence state of insulating 4f compounds detected by $L_3$ X-ray absorption. Physical Review B, 35(2):806-812.
\bibitem{Bianconi Kotani1988}
Bianconi, A., Kotani, A., Okada, K., Giorgi, R., Gargano, A., Marcelli, A., and Miyahara, 
T. (1988). Many-body effects in praesodymium core-level spectroscopies of $PrO_2$. Physical Review B, 38(5):3433-3437.
\bibitem{BianconiClozza1989}
Bianconi, A., Clozza, A., Murata, T., Matsukawa, T., Miyahara, T., Kotani, A., Nakai, S., and Mitsuishi, T. (1989). 
Many body effects in Ce 3p XAS and Ce 3p XPS of Ce$O_2$. Physica B: Condensed Matter, 158(1-3):389-391.
\bibitem{BianconiMiyahara1989}
Bianconi, A., Miyahara, T., Kotani, A., Kitajima, Y., Yokoyama, T., Kuroda, H., Funabashi, M., Arai, H., and Ohta, T. (1989). 
Correlation satellites in deep metal 3p core x-ray photoemission of tetravalent oxides M$O_2$ (M=Ce,Pr,Tb,Hf) and of $LaF_3$. Physical Review B, 39(5):3380-3385.

\bibitem{BianconiRudolf1987} 
Bianconi, A., Congiu-Castellano, A., De Santis, M., Rudolf, P., Lagarde, P., Flank, A. M., and Marcelli, A. (1987)
 $L_{2,3}$  XANES of the high  superconductor $YBa_2Cu_3O_7$ with variable oxygen content. Solid State Communications, 63(11):1009-1013.
 
\bibitem{BianconiPolitis1987}
Bianconi, A., Castellano, A., Santis, M., Politis, C., Marcelli, A., Mobilio, S., and Savoia, A. (1987). 
Lack of delocalized Cu p states at the Fermi level in the high-$T_c$ superconductor $YBa_2Cu_3O_7$ 
by  XANES spectroscopy.  Zeitschrift für Physik B Condensed Matter, 67(3):307-312.

 \bibitem{BianconiDelogu1987}
 Bianconi, A., Congiu-Castellano, A., De Santis, M., Delogu, Gargano, and Giorgi (1987). 
Localization of Cu 3d levels in the high $T_c$ superconductor $YBa_2Cu_3O_7$ by 
Cu $2p$ x-ray photoelectron spectroscopy. Solid State Communications, 63(12):1135-1139.
  
\bibitem{BianconiClozza1987}
Bianconi, A., Clozza, A., Congiu Castellano, A., Della Longa, S., De-Santis, M., Di Cicco, A., Garg, K., Delogu, P., Gargano, A., Giorgi, R., Lagarde, P., Flank, A. M., 
and Marcelli, A. (1987). Experimental evidence of itinerant Cu($3d^9$)-Oxygen-hole many body configuration in the high-$T_c$ superconductor $YBa_2Cu_3O_7$.
 International Journal of Modern Physics B (IJMPB), 1(3-4):853-862.
 
 \bibitem{BianconiClozzaDiCicco1987}
 Bianconi, A., Clozza, A., Congiu Castellano, A., Della Longa, S., De Santis, M., 
Di Cicco, A., Garg, K., Delogu, P., Gargano, A., Giorgi, R., Lagarde, P., Flank, A. M., and Marcelli, A. (1987). 
Cu 3$d^9$ - Ligand Hole Configuration In $YBa_2 Cu_3 O_{~7}$ by X-ray spectroscopies. Le Journal de Physique Colloques, 48(C9):C9-1179-C9-1184.

\bibitem{KanamoriKotani1988}
Kanamori, J. and Kotani, A. (1988). 
Core-Level spectroscopy in condensed systems proceedings of the tenth Taniguchi international symposium, Kashikojima, Japan, october 19-23, 1987.
   
\cite{BianconiBudnickChamberland1988}
\bibitem{BianconiBudnickChamberland1988}
Bianconi, A., Budnick, J., Chamberland, B., Clozza, A., Dartyge, E., Demazeau, G., De Santis, M., Flank, A. M., Fontaine, A., Jegoudez, J., Lagarde, P., Lynds, L. L., Michel, C., Otter, F. A., Tolentino, H., Raveau, B., and Revcolevschi, A. (1988). $3d^9L$ states induced by doping in $La_{1.85}Sr_{0.15}CuO_4$ and in magnetic and non magnetic $M_1Ba_2Cu_3O_7$ (M = Gd, Ho and Y). Physica C: Superconductivity, 153-155:113-114.

 \bibitem{BianconiBudnickDemazeau1988}
 Bianconi, A., Budnick, J., Demazeau, G., Flank, A. M., Fontaine, A., Lagarde, P., Jegoudez, J., Revcolevski, A., Marcelli, A., and Verdaguer, M. (1988). 
Cu $L_3$ X-ray absorption of formally trivalent Cu compounds. Physica C: Superconductivity, 153-155:117-118.

\bibitem{Grioni1987}
Grioni, M., Fuggle, J. C., Weijs, Goedkoop, J. B., Rossi, G., Schaefers, F., Fink, J., and Nuecker, N. (1987).
High energy spectroscopies of Cu compounds including new superconductors. Le Journal de Physique Colloques, 48(C9):C9-1189-C9-1192.
  
\bibitem{Nuecker1987}
Nuecker, N., Fink, J., Renker, B., Ewert, D., Politis, C., Weijs, P. J. W., and Fuggle, J. C. (1987). 
Experimental electronic structure studies of $La_{2-x }$$Sr_x$$CuO_4$.  Zeitschrift für Physik B Condensed Matter, 67(1):9-14.

\bibitem{Nuecker1988}
Nuecker, N., Fink, J., Fuggle, J. C., Durham, P. J., and Temmerman, W. M. (1988). 
Evidence for holes on oxygen sites in the high- Tc superconductors $La_{2-x}Sr_xCuO_4$ and $YBa_2Cu_3O_{7-y}$. Physical Review B, 37(10):5158-5163.

\bibitem{BianconiDesantisFlank1988}
Bianconi, A., De Santis, M., Flank, A., Fontaine, A., Lagarde, P., Marcelli, A., Katayama-Yoshida, H., and Kotani, A. (1988). 
Determination of the symmetry of the 3$d^9$L states by polarized Cu $L_3$ XAS spectra of single crystal $YBa_2Cu_3O_{6.9}$. Physica C: Superconductivity, 153-155:1760-1761.
  
\bibitem{BianconiDeSantisDiCicco1988}
Bianconi, A., De Santis, M., Di Cicco, A., Flank, A., Fontaine, A., Lagarde, P., Katayama-Yoshida, H., Kotani, A., and Marcelli, A. (1988). 
Symmetry of the $3d^9$ ligand hole induced by doping in $YBa_2Cu_3O_{7-y}$ Physical Review B, 38(10):7196-7199.

\bibitem{BianconiDeSantisDiCicco1989}
Bianconi, A., De Santis, M., Dicicco, A., Flank, A., and Lagarde, P. (1989).
Symmetry of the itinerant $3d^9$ ligand hole induced by doping in high $T_c$ superconductors by polarized Cu $L_3$ XAS. Physica B: Condensed Matter, 158(1-3):443-445.

\bibitem{Flank1990}
Flank, A. M., Lagarde, P., Bianconi, A., Castrucci, P., Fabrizi, A., Pompa, M., Katayama-Yoshida, H., and Calestani, G. (1990). 
High resolution polarized Cu $L_3$ XAS of high temperature superconductors: correlations between the symmetry of d holes and the critical temperature. Physica Scripta, 41(6):901-904.

\bibitem{Seino1990}
Seino, Y., Kotani, A., and Bianconi, A. (1990). Effect of rhombic distortion on the polarized X-ray absorption spectra in high $T_c$ superconductors. Journal of the Physical Society of Japan, 59(3):815-818.

\bibitem{BianconiMissoriOyanagi1995}
Bianconi, A., Missori, M., Oyanagi, H., Yamaguchi, H., Nishiara, Y., Della Longa, S., (1995) The measurement of the polaron size in the metallic phase of cuprate superconductors. Europhysics Letters (EPL) 31 (7):411-415.
% http://iopscience.iop.org/0295-5075/31/7/012
\bibitem{Li}
Li, C., Pompa, M., Congiu-Castellano, A., Della-Longa, S., and Bianconi, A. (1991). 
One-electron excitations and shake up satellites in Cu K-edge x-ray absorption near edge structure (XANES) of $La_2CuO_4$ by full multiple scattering analysis in real space. Physica C: Superconductivity, 175(3-4):369-380.
\bibitem{BianconiLiCampanella1991} 
Bianconi, A., Li, C., Campanella, F., Longa, S. D., Pettiti, I., Pompa, M., Turtu, S., and Udron, D. (1991). 
Cu K-edge polarized X-ray-absorption near-edge structure of $Bi_2CaSr_2Cu_2O_8$. Physical Review B, 44(9):4560-4569.
\bibitem{Pompa 1991}
Pompa, M., Li, C., Bianconi, A., Congiu-Castellano, A., della Longa, S., Flank, A. M., Lagarde, P., and Udron, D. (1991). 
Full multiple scattering analysis of linearly polarized Cu $L_3$-edge XANES of $La_2CuO_4$. Physica C: Superconductivity, 184(1-3):51-64.
 \bibitem{BianconiDellaLongaLi1991}
Bianconi, A., Della Longa, S., Li, C., Pompa, M., Congiu-Castellano, A., Udron, D., Flank, A. M., and Lagarde, P. (1991). 
Linearly polarized Cu $L_3$-edge x-ray-absorption near-edge structure of $Bi_2CaSr_2Cu_2O_8$ . Physical Review B, 44(18):10126-10138.
\bibitem{Pellegrin1993}
Pellegrin, E., Nuecker, N., Fink, J., Molodtsov, S. L., Gutiérrez, A., Navas, E., Strebel, O., Hu, Z., Domke, M., Kaindl, G., Uchida, S., Nakamura, Y., Markl, J., Klauda, M., Saemann-Ischenko, G., Krol, A., Peng, J. L., Li, Z. Y., and Greene, R. L. (1993). Orbital character of states at the fermi level in $La_{2-x}Sr_xCuO_{4}$  and $R_{2-x}Ce_xCuO_4 $ (R=Nd,Sm). Physical Review B, 47(6):3354-3367.
\bibitem{Pellegrin1995} Pellegrin, E., Chen, C. T., Zaanen, J., and Fink, J. (1995). Symmetry of hole states in $La_{2-x}Sr_xCuO_{4 +y}$ and   $La_{2-x}Sr_xNiO_{4 +y}$. Physica B: Condensed Matter, 208-209:487-490.
\bibitem{Pellegrin1996} Pellegrin, E., Fink, J., Chen, C. T., Xiong, Q., Lin, Q. M., and Chu, C. W. (1996). Experimental hole densities in $HgBa_2Ca_{n-1}Cu_nO{2n+2}$ compounds from near-edge x-ray-absorption spectroscopy. Physical Review B, 53(5):2767-2772.
\bibitem{Merz1997}
Merz, M., Nuecker, N., Pellegrin, E., Schweiss, P., Schuppler, S., Kielwein, M., Knupfer, M., Golden, M. S., Fink, J., Chen, C. T., Chakarian, V., Idzerda, Y. U., and Erb, A. (1997). 
X-ray absorption spectroscopy of detwinned $Pr_xY_{1-x}Ba_2Cu_3O_{7-y}$ single crystals: electronic structure and hole distribution. Physical Review B, 55(14):9160.
\bibitem{Valletta1997}
Valletta, A., Bianconi, A., Perali, A., Saini, N. L. (1997). Electronic and superconducting properties of a superlattice of quantum stripes at the atomic limit. Zeitschrift fur Physik B Condensed Matter 104 (4): 707-713.
\bibitem{bianconi97}
Bianconi, A., Valletta, A., Perali, A., and Saini, N. L. (1997). High Tc superconductivity in a superlattice of quantum stripes. Solid State Communications, 102(5):369-374.
\bibitem{Perali2012}
Perali, A., Innocenti, D., Valletta, A., Bianconi, A., (2012). Anomalous isotope effect near a 2.5 Lifshitz transition in a multi-band multi-condensate superconductor made of a superlattice of stripes. Superconductor Science and Technology 25 (12):124002.
\bibitem{h3s1}
Bianconi, A. and Jarlborg, T. (2015). Superconductivity above the lowest earth temperature in pressurized sulfur hydride. EPL (Europhysics Letters), 112(3):37001+.
\bibitem{h3s2}
Bianconi, A. and Jarlborg, T. (2015). Lifshitz transitions and zero point lattice fluctuations in sulfur hydride showing near room temperature superconductivity. Nov. Supercond. Mater., 1(1):37-49.
\bibitem{h3s3}
Jarlborg, T. and Bianconi, A. (2016). Breakdown of the migdal approximation at lifshitz transitions with giant zero-point motion in the H3S superconductor. Scientific Reports, 6:24816 
\bibitem{h3s4}
Bussmann-Holder, A., Kohler,J., Whangbo, M.-H., Bianconi,A., Simon, A. (2016). High temperature superconductivity in sulfur hydride under ultrahigh pressure: 
A complex superconducting phase beyond conventional BCS.
Nov. Supercond. Mater.  2: 37-42 DOI 10.1515/nsm-2016-0004




 \end{thebibliography}

\documentclass[twocolumn,showpacs,preprintnumbers,amsmath,amssymb]{revtex4}
\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}

\begin{document}

\title{Many body quantum physics in XANES of highly correlated materials, mixed valence oxides and high temperature superconductors}

\author{Antonio Bianconi$^{1,2,3}$}

\affiliation{$^1$ RICMASS, Rome International Center for Materials Science Superstripes, Via dei Sabelli 119A, 00185 Rome, Italy}
\affiliation{$^2$ Solid State and Nanosystems Physics,  National Research Nuclear University, MEPhI, Moscow Engineering Physics Institute,
       Kashirskoye sh. 31, Moscow 115409, Russia}
\affiliation{$^3$ CNR, Institute of Crystallography, Consiglio Nazionale delle Ricerche, via Salaria, 00015 Monterotondo, Italy}


\begin{abstract}

The x-ray absorption near edge structure (XANES), developed in these last 40 years using synchrotron radiation, is a unique tool probing electronic correlations in complex systems via quantum many body final state effects.  Multi electron excitations have been observed first in the sixties in x-ray absorption spectra of atoms and later in molecules and solids. The applications of XANES many body final states to probe unique features of electronic correlation in heavy fermions, mixed valence systems, mixed valence oxides and high temperature superconductors are discussed.

\end{abstract}

\pacs{78.70.Dm, 71.10.-w, 74.72.-h,71.27.+a,71.28.+d, 75.20.Hr, 75.30.Mb }

\maketitle

\section{Introduction.}

The origin of quantum many body physics is usually associated with the 1926 publication of Heisenberg 
paper \cite{Heise} where he provided the theoretical interpretation of the energy splitting of the spectral
 lines, called ortho- and para- helium, of helium atomic gas absorption spectra.
In this work Heisenberg assumed that the motion of the two electrons in helium is correlated considering it as
the simplest many body electronic system made of only two electrons.
Therefore he described a quantum resonance between two electronic configurations of the two electrons.
In this work he introduced the exchange interaction giving the energy splitting between symmetric 
and antisymmetric wave-functions.
Following the Fermi (1926) article  \cite{fermi} on the Fermi statistics for the electron gas in metals, Dirac (1926) \cite{Dirac}
introduced the distinction of the Fermi versus Bose statistics for fermions  versus bosons characterized by
antisymmetric versus symmetric wave-function respectively. The exchange interaction between 
bosons is attractive, giving at T=0 a quantum Bose condensate,
while the exchange interaction is repulsive for fermions giving at T=0 the degenerate Fermi gas.
Wentzel (1927) \cite{Wentzel} proposed a non trivial extension of the Heisenberg (1926) theory \cite{Heise} to explain the Auger effect.

\section{many body effects in atomic physics}

Majorana (1931) \cite{majorana} extended the Wentzel theory focusing on the selection rules for the non-radiative decay of two electron excitations observed in atomic absorption spectra (Foote \textit{et al.} 1925) \cite{Foote}, (Shenstone, 1931) \cite{Shenstone}.
Following the experiment of Beutler (1935)  \cite{Beutler} on vacuum ultraviolet absorption spectra of rare gases where  broad lines 
due to multi-electron excitations were observed beyond the ionization potential, Ugo Fano (Fano, 1935) \cite{Fano} developed the theory of configuration interaction between discrete and continuum channels with the prediction of the asymmetric line-shape of absorption spectrum of two electron excitations. 

The Fano theory was tested by  the experimental line-shape of shape resonances in nuclear scattering experiments due to configuration interaction between open and closed scattering channels (Blatt \& Weisskopf, 1952)   \cite{Blatt1}, (Blatt \textit{et al.},1953) \cite{Blatt2}, (Feshbach \textit{et al.}, 1954) \cite{Fes1} . This theory was extended to many different scattering channels in many body systems by Feshbach (1958) \cite{Fes2}. The Fano-Feshbach resonances are today of high scientific interest in many different fields and in particular for ultracold gases and high-$T_c$ superconductivity (Vittorini-Orgeas \&  Bianconi, 2009)  \cite{Vittorini}.

Ugo Fano at National Bureau of Standard in 1959 promoted synchrotron radiation research using a 180 MeV synchrotron to measure high resolution soft x-ray absorption spectroscopy of atoms. He proposed to investigate the line-shape of two electron excitations, and published a second extended version  of his 1935 paper in Physical Review (Fano 1961) \cite{Fano2}. 
The experiments  (Madden \&  Codling, 1963) \cite{Madden1}, (Madden \&  Codling, 1965) \cite{Madden2}  provided a clear experimental confirmation of the asymmetric Fano line-shape   
for two electron excitations in helium, neon and argon where the interference of a discrete autoionized state with a continuum gives rise to characteristically asymmetric peaks in soft x-ray absorption spectra (Cooper \textit{et al.}, 1963) \cite{Cooper}. 

The Fano-Feshbach resonance in many body configuration interaction between discrete and continuum states is a fundamental quantum process first observed in x-ray absorption spectra but it was soon applied in condensed matter physics to describe magnetic impurities embedded in a metallic host, in the so-called Anderson Impurity Model  (Anderson, 1961) \cite{Anderson} and in the prediction of the characteristic change in electrical resistivity with temperature known as Kondo effect  (Kondo, 1964) 
\cite{Kondo} which is due to the scattering of conduction electrons by magnetic impurities. Today it is very popular in a variety systems ranging from heavy fermions to Kondo insulators. 

The investigations of two-electron excitations have been extended to different excitations  at high energies in helium Dhez \&  Ederer 1973 \cite{dhez}
and in kripton  (Bernieri  \& Burattini 1987) \cite{bernieri}. The two-electron excitations have been found later in solids, as for example in solid neon (Soldatov \textit{et al.}, 1993) \cite{soldatov} in V, Cr, and Mn $3d^0$ compounds with tetrahedral coordination
 (Bianconi \& Garcia \& Benfatto \textit{et al.}, 1991)  \cite{bianconi1991a}
   in  $SiCl_3$ and $SiF_3$ (Di Cicco \textit{et al.},1992) \cite{dicicco}, in hydrated calcium (D Angelo \textit{et al.}, 2004) \cite{dangelo} and in rare earth alloys  (Chaboy,1990) \cite{chaboy1990}, (Chaboy,1994) \cite{chaboy1994}.

\section{Many body effects in condensed matter}

In 1970 the interest of the synchrotron radiation research in novel facilities using small synchrotrons in Madison (USA) and Tokyo (Japan) and large synchrotrons in Hamburg (Germany)  and and Frascati (Italy) shifted from atomic physics to condensed matter.  New many body theories predicted many body final state effects at the absorption threshold of metals in the soft x-ray absorption spectra  (Roulet \textit{et al.}, 1969) \cite{roulet} (Nozieres \& De Dominicis, 1969) \cite{nozieres} (Schotte \&   Schotte, 1969),  \cite{shotte} (Mahan, 1971) \cite{mahan}  (Hedin \textit{et al.}, 1971) \cite{heidin} (Combescot \&  Nozieres, 1971) \cite{combescot}. 
Similar many body effects predicted in x-ray photoelectron spectroscopy (XPS) (Doniach \&  Sunjic, 1970)  \cite{doniach}. for metals have been supported  by the measurements of asymmetric line-shape of core level spectra (Antonangeli \textit{et al.}, 1977) 
 \cite{antonangeli}. . 
Many body final states effects related with the core hole excitation have been observed in the x-ray absorption spectra of palladium hydrides  (Benfatto \textit{et al.}, 1983)   \cite{benfatto1983} and  in photoemission cross-section  near the carbon K shell  threshold, many body plasmon excitations have been shown to play a relevant role (Bachrach \& Bianconi, 1982) 
\cite{bachrach1982}.
However the many body effects in x-ray absorption spectra of metals give only subtle changes of the line-shape of absorption threshold which are difficult to be observed.

\section{Shape resonances in molecules and condensed matter}

The interest of the scientific community has been focussed 
mainly to understand one-electron final states effects due 
to the role of density of states (DOS)  in the conduction band as
 shown for example in the edge region of the $L_2,$$_3$ absorption spectrum of aluminum (Balzarotti \& Bianconi  \& Burattini 1974)  \cite{balz1974} measured at the facility Solidi Roma at the 1 GeV synchrotron in Frascati, Italy.
Synchrotron radiation spectroscopy using high energy and large accelerators (3.5 GeV SPEAR in Stanford, 7.4 GeV Desy in Hamburg) focused on the weak features in the high energy range called x-ray absorption fine structure (EXAFS). The theory and experiments on weak EXAFS oscillation in a variety of complex systems evolved rapidly from 1970 to 1980. On the contrary the theoretical interpretation of the low energy part of the x-ray absorption spectra giving strong absorption absorption peaks in the continuum of solids, both crystals and glasses (Balzarotti \& Bianconi \& Burattini \textit{et al.}, 1974)  \cite{balza1974} 
remained obscure. These peaks were qualitatively assigned to transitions to unoccupied DOS of the crystal or atomic final states. The physics was clarified by the observation of the shape resonance in the continuum of K-edge absorption of the nitrogen molecule
 (Bianconi \& Petersen \& Brown \textit{et al.}, 1978) \cite{BianconiPetersen1978}. 
 
 The shape resonance is a short living final state made of a quasi stationary state, formed by a multiple scattering resonance of the photoelectron, degenerate with the continuum. Its spatial localization in a nanoscale cluster is determined by the high scattering amplitude of the low kinetic energy, 10-50 eV, of the photoelectron from neighbor atoms. The very short life time, of the order of $10^{-15} s$, is determined by the core hole life time and the shape resonance decay into the continuum.  The shape resonances are related with the effects of electronic correlation in molecules 
(Langhoff \& Davidson 1974) \cite{Langhoff1974}, 
(Broad \& Reinhardt 1976) \cite{Broad1976}, 
(Dehmer \& Dill, 1975) \cite{Dehmer1975}, 
(Dehmer \& Dill, 1976) \cite{Dehmer1976}, 
(Dehmer \textit{et al.}, 1979) \cite{Dehmer1979},
(Loomba \textit{et al.}, 1981) \cite{Loomba1981}.  In a set of papers (Bianconi \& Doniach \& Lublin 1978) \cite{BianconiDoniach1978} 
 (Bianconi, 1979) \cite{Bianconi1979}
  (Bianconi, 1980)  \cite{Bianconi1980}
 the acronym  XANES (x-ray absorption near edge structure) was coined to indicate the spectral features in the absorption spectra of condensed matter, including biological systems, chemical compounds, liquids, amorphous systems and glasses due to shape resonances for the photoelectron which in the final state is confined for a short time in a nanoscale cluster centered on the selected absorption atom. The XANES energy range, extending over 50-100 eV beyond the absorption edge, is defined where the photoelectron wavelength is larger that the interatomic distance. Here the full multiple scattering theory is needed to predict the observed multiple-scattering resonances and it was shown that the XANES variation due subtle structural changes in Fe II and Fe III hexacyanide complexes can be predicted and understood
 (Bianconi \& Dell'Ariccia \& Durham \textit{et al.}, 1982) \cite{Bianconi1982}. 

The XANES data analysis (Della Longa \textit{et al.}, 1995)   \cite{DellaLonga1995}
using the full multiple scattering theory provides today the 
coordination geometry and higher order structural correlation function in the liquid phase 
(Garcia \&  Bianconi \&  Benfatto, \textit{et al.}, 1986)  \cite{garciaBianconi1986}, 
(Garcia \&  Benfatto \&  Natoli \textit{et al.}, 1986) \cite{garcia1986}
 and in a large variety of disordered systems.

\section{Mixed valence systems and heavy fermions}

 The search for many body final states in XANES focused on  valence fluctuating materials like Yb inter-metallics  
(Rao  \textit{et al.}, \cite{Rao1980}  1980) and
TmSe (Launois \textit{et al.},1980)  \cite{Launois1980} 
(Bianconi \& Modesti \& Campagna \textit{et al.},1981)    \cite{BianconiModesti1981} 
and $CePd_3$ (Sarode \textit{et al.}, 1981)  \cite{Sarode1981} 
  were it was shown that the quantitative  mixed valence state could be extracted from the data.
  
After the discovery of superconductivity in the first  heavy fermion system $CeCu_2Si_2$ (Steglich \textit{et al.}, 1979) \cite{Steglich1979}  few groups started to look for experimental evidence of strong electronic correlation in intermetallics. Heavy fermions are at the edge between superconductivity and magnetism. In heavy fermions at the chemical potential a narrow f band, with a large effective mass, coexists with a wide d band 
(Jarlborg \textit{et al.}, 1983). \cite{Jarlborg1983} \ These materials show strong Pauli paramagnetism and electronic complexity (Lieke \textit{et al.},1982) \cite{Lieke1982} 
and their electronic properties emerge from multi-electron electronic configurations.
The breakdown of single  electron approximation in the XANES of heavy fermions like $CeCu_2Si_2$ 
(Bianconi \& Campagna \& Stizza \textit{et al.}, 1981) \cite{BianconiCampagna1981}  and 
$EuPd_2Si_2$ \cite{Padalia1981}  has been found and evidence for the competition between two many body electronic configurations was clearly detected by XANES. 

   The intermediate valence probed by core-level spectroscopy  has  been investigated  in light  Ce vs  heavy Tm rare earths. Unique information on electronic correlation, localization and local-environment effects have been obtained by L-edge XANES
 (Bianconi \& Campagna \& Stizza  \textit{et al.}, 1982).   \cite{BianconiCampagna1982}  
The changes of many body electronic configurations at the Fermi level at the phase transition from $\alpha$ to $\gamma$ phase of cerium have been observed at the Ce $L_3$ absorption edge 
(Lengeler \textit{et al.}, 1983) \cite{Lengeler1983}. 
The effects of dilution and chemical pressure on mixed valence cerium-based systems like $CeNi_5$ have been studied by (Raaen \& Parks, 1983) \cite{Raaen1983} (Parks  \textit{et al.}, 1983)  \cite{Parks1983}.
The role of ocalization and mixing of many body configurations in intermetallic compounds $RPd_3$ (R= La, Ce, Pr, Nd, Sm) have been measured by
 (Bianconi \& Marcelli  \& Davoli \textit{et al.}, 1984)    \cite{BianconiMarcelli1984} 
 and (Marcelli \textit{et al.},1984)  \cite{Marcelli1984} 
 (Marcelli, \textit{et al.}, 1985)   \cite{Marcelli1985},
(Bianconi \& Marcelli  \textit{et al.}, 1985) \cite{BianconiMarcelli1985} 
  by $L_3$ and $L_2$-XANES spectra. The many body final states 
  in inner shell photoemission and photoabsorption spectra of La and Ce 
  strongly correlated electronic compounds have been interpreted theoretically using  the Anderson impurity model 
  (Kotani \& Jo \& Okada  \textit{et al.}, 1987) \cite{KotaniJo1987},   
  (Kotani \& Okada \& Jo \textit{et al.}, 1987). \cite{KotaniOkada1987}

\section{Strongly correlated systems: charge transfer Mott insulators and insulator to metal transitions}

While in the x-ray absorption near edge spectroscopy a very high energy photon (in the KeV energy range) is absorbed by a deep core level, the excited electron in the final state occupies the lowest unoccupied electronic states at the chemical potential therefore many body  
electronic configurations ib correlated electronic systems determine the XANES final states. This feature was used to probe the variation of many body electronic configurations at  the metal-insulator transition in $V_2O_3$ (Bianconi \& Natoli,1978)  \cite{BianconiNatoli1978}    
and in $VO_2$ (Bianconi,1982)  \cite{Bianconi1982}  measuring the high resolution vanadium K-photoabsorption spectrum. 
 In 1985 the theory of correlation gaps in the electronic structure of transition-metal compounds 
 was presented (Zaanen \textit{et al.},1985)  \cite{Zaanen1985} showing the existence of different class of Mott insulators involving both the metal and ligand orbitals. 
 Davoli \textit{et al.} in 1986  \cite{Davoli1986}   provided direct evidence for the correlation gap between $3d^8$  and $3d^9L$ multi-electron configurations (where L indicated the hole in the ligand orbital) in the x-ray-absorption near-edge structure of NiO 
  using oxygen K-edge XANES in agreement with predictions of (Zaanen \textit{et al.}, 1985) \cite{Zaanen1985}.
  Many body configurations in the x-ray absorption spectra of charge transfer gap insulators have been measured in other correlated systems like nickel dihalides   (van der Laan  \textit{et al.}, 1986) \cite{van der Laan1986} in agreement with the teory  (Zaanen \textit{et al.}, 1986)  \cite{Zaanen1986},  uranium 5f oxides and glasses, (Petiau, \textit{et al.}, 1986) \cite{Petiau1986},  
  and 4f compounds (Bianconi \& Marcelli \& Dexpert \textit{et al.}, 1987) \cite{BianconiMarcelliDexpert1987}  like $PrO_2$  (Bianconi \& Kotani \textit{et al.}, 1988) \cite{Bianconi Kotani1988}  and in in $Ce(SO_4)_2$ (Bianconi \& Marcelli  \& Tomellini \textit{et al.}, 1985) \cite{BianconiMarcelli1985}. A full information on ground state many body configurations  of $CeO_2$, (M$O_2$ (M=Ce,Pr,Tb,Hf) and of $LaF_3$ has been  obtained by joint Ce 3p XANES and  deep metal 3p core x-ray photoemission spectroscopy XPS using hard x-rays spectroscopy of tetravalent oxides (Bianconi \& Clozza \textit{et al.}, 1989)  \cite{BianconiClozza1989},  (Bianconi \& Miyahara \& Kotani \textit{et al.}, 1989) \cite{BianconiMiyahara1989}.

\section{high temperature superconductors}

The discovery in 1986 of high temperature superconductivity in copper oxides triggered a large interest on the nature of new metallic states induced by chemical doping of strongly correlated layered copper oxides.
The strongly correlated many body state of the copper oxide plane was studied by Cu $L_{2,3}$-XANES (Bianconi \& Congiu-Castellano  \& De-Santis \& Rudolf, \textit{et al.}, 1987)  \cite{BianconiRudolf1987}, by Cu $K$-edge XANES (Bianconi \& Castellano \& De Santis  \& Politis \textit{et al.}, 1987)  \cite{BianconiPolitis1987}  and by Cu 2p x-ray photoelectron spectroscopy (Bianconi \& Congiu-Castellano \&  De-Santis \&  Delogu \textit{et al.}, 1987) \cite{BianconiDelogu1987}.
These experiments provided the first direct measure of the Mott charge transfer gap between the many body configurations Cu$3d^9$-O$2p^{10}$  (called $3d^9$) and Cu$3d^{10}$-O$2p^5$ (called $3d^{10}$L where L is the oxygen hole) in the parent compounds which was measured to be about 2 eV. These systems can be turned from insulator to metal by variable oxygen interstitials content y like in the high $T_c$ superconductor $YBa_2Cu_3O_{6+y}$. The variation of the many body final states in XANES with variable oxygen content  provided the first direct experimental evidence that the itinerant Cu$3d^9$L (where L is the hole on oxygen orbital) many body configuration are created at the Fermi level in the correlation gap by doping  (Bianconi \& Congiu-Castellano  \& De-Santis \& Rudolf, \textit{et al.}, 1987) \cite{BianconiRudolf1987},  
(Bianconi \& Clozza \& Congiu-Castellano \& Della-Longa \textit{et al.}, 1987) \cite{BianconiClozza1987}, (Bianconi \& Clozza, \& Congiu-Castellano, \& Della Longa, \&  De Santis, \&  Di Cicco, \textit{et al.}, 1987)  \cite{BianconiClozzaDiCicco1987} and these results were presented at the Taniguchi symposium, Kashikojima, Japan, october 19-23, 1987.(Kanamori \& Kotani, 1988)   \cite{KanamoriKotani1988}. In 1988  these results have been confirmed by similar findings in other cuprates (Bianconi \& Budnick \& Chamberland \textit{et al.},  1988)   \cite{BianconiBudnickChamberland1988} and in formally trivalent Cu compounds (Bianconi \& Budnick \& Demazeau \textit{et al.}, 1988)  \cite{BianconiBudnickDemazeau1988}.
While other authors presented different interpretations of the XANES data in 1987 (Grioni \textit{et al.}, 1987)  \cite{Grioni1987}, (Nuecker \textit{et al.}, 1987)  \cite{Nuecker1987}, the oxygen K-edge  absorption  (Nuecker  \textit{et al.}, 1988) \cite{Nuecker1988} provided a compelling evidence for the formation of  Cu$3d^9$L  many body states induced by doping. 
 
 In 1988 using polarized XANES spectroscopy of single crystals it was shown that there are two coexisting different many body configurations:  Cu$3d^9$L ($b_1$), where the holes on oxygen orbital have fully planar symmetry L($b_1$)  and and Cu$3d^9$L($a_1$) where the oxygen holes have partially out-of-plane symmetry  L($a_1$)  (Bianconi \& Desantis \& Flank \textit{et al.}, 1988)  \cite{BianconiDesantisFlank1988}, (Bianconi \& De Santis \& Di Cicco \textit{et al.}, 1988) \cite{BianconiDeSantisDiCicco1988}, 
 (Bianconi \&  De Santis \& Di Cicco \textit{et al.}, 1989) \cite{BianconiDeSantisDiCicco1989}. 
  
The presence of two states with different orbital symmetry is relevant for superconductivity (Flank \textit{et al.}, 1990)   \cite{Flank1990}.
 The presence of a second electronic component Cu$3d^9$L($a_1$) was shown to be associated with the rhombic distortion on the $CuO_4$ square plane (Seino  \textit{et al.},1990) \cite{Seino1990} related to charge density wave onset and polaron formation (Bianconi  \&   Missori \&  Oyanagi, \textit{et al.}, 1995)  \cite{BianconiMissoriOyanagi1995}. 
 
 Full multiple scattering analysis of polarized K-edge  (Li  \textit{et al.}, 1991) \cite{Li},  (Bianconi \&  Li \& Campanella, \textit{et al.}, 1991)  \cite{BianconiLiCampanella1991},   
 and  Cu $L_3$-edge (Pompa \textit{et al.}, 1991)  \cite{Pompa 1991},
 (Bianconi \&  Della Longa \& Li \textit{et al.}, 1991) \cite{BianconiDellaLongaLi1991} 
 XANES of $La_2CuO_4$ and $Bi_2CaSr_2Cu_2O_{8+y}$ have clearly shown that many body final states configurations give a splitting of the strongest peak in the Cu K-edge XANES.
  The presence of two set of states at the Fermi level with different orbital character has been confirmed by many experiments at O 1s and Cu 2p edges in  $La_{2-x}Sr_xCuO_{4+y}$, $La_{2-x}Sr_xNiO_{4+y}$, $(Y,R)Ba_2Cu_3O_7$  
 (Pellegrin \textit{et al.}, 1993)  \cite{Pellegrin1993},
  $Bi_2CaSr_2Cu_2O_{8+y}$ 
 (Pellegrin \textit{et al.}, 1995)  \cite{Pellegrin1995} 
 and $HgBa_2Ca_{n-1}Cu_nO_{2n+2+y}$ 
 (Pellegrin \textit{et al.}, 1996)   \cite{Pellegrin1996},  $Pr_xY_{1-x}Ba_2Cu_3O_{7-y}$ 
 (Merz, \textit{et al.}, 1997) \cite{Merz1997}.  Today the  two electronic many body components seen by XANES provide the experimental base for the recent development of the theory of multigap superconductivity both in cuprates  
 (Valletta \textit{et al.}, 1997,Bianconi textit{et al.}, 1997) \cite{Valletta1997,bianconi97}
 (Perali \textit{et al.}, 2012) \cite{Perali2012} and other high temperature superconductors including pressurized sulfur hydrides \cite{h3s1,h3s2,h3s3.h3s4}.

\end{document}